OpenStudy (anonymous):

Find the maximum positive x such that a) sin x=0 b) cos 2x-1/2=0 c) sin (x^2 + x + 4)=0

OpenStudy (anonymous):

\(\sin x=0\) for \(x=2n\pi\), for \(n=0,\pm1,\pm2,...\). There's no maximum for the first one, unless you restrict the domain. Is the second one \(\cos2x-\dfrac{1}{2}=0\) or \(\cos\left(2x-\dfrac{1}{2}\right)=0\) ? The third is a lot like the first, in that \[\sin(x^2+x+4)=0\] for \[x^2+x+4=2n\pi,~\text{for }n=0,\pm1,\pm2,...\]